3.10 If U, V are ideals of R, let U + V = {u + v | u ∈ U, v ∈ V }.

Prove that U + V is also an ideal.


1
Expert's answer
2022-06-28T15:30:16-0400

Let "u + v \in U + V" and "r \in R". Since "U" is an ideal, "\{ru, ur\} \subset R"; since "V" is an ideal, "\{rv, vr\} \subset R". Hence, using left and right distributivity of the ring "R",

"r(u + v) = ru + rv \in R, \\
(u + v)r = ur + vr \in R"

(the sums are elements of "R" because the ring is stable under addition). This proves that "U + V" is both a left and right ideal of "R". Therefore, "U + V" is an ideal.